Two particles P and Q move towards each other along a straight line MN, 51 meters long. P starts from M with velocity 5ms-1 and constant acceleration of 1ms-2. Q starts from N at the same time with velocity 6ms-1 and at a constant acceleration of 3ms-2.


Find the time when the :

A) particles are 30 metres apart

B)Particles meet

C) velocity of P is ¾ of the velocity of Q


1
Expert's answer
2020-07-13T11:43:57-0400

(a)


"s_1=v_{01}t+\frac{a_1t^2}{2}" and "s_2=v_{02}t+\frac{a_2t^2}{2}"


"51-(s_2+s_1)=30 \to s_2+s_1=21"


"v_{01}t+\frac{a_1t^2}{2}+v_{02}t+\frac{a_2t^2}{2}=21 \to 5t+\frac{t^2}{2}+6t+\frac{3t^2}{2}=21 \to"


"t=1.5s"


(b)


"5t+\frac{t^2}{2}+6t+\frac{3t^2}{2}=51 \to t=3s"


(c)


"v_1=v_{01}+a_1t=5+t" and "v_2=v_{02}+a_2t=6+3t"


"v_1=\frac{3}{4}v_2\to 5+t=4.5+2.25t\to t=0.4s"